# TorchCP: A Python Library for Conformal Prediction

### v3

@@ Abstract
Conformal prediction (CP) is a robust statistical framework that generates prediction intervals or sets with guaranteed coverage probability, addressing the challenge of quantifying predictive uncertainty in deep learning. Despite advancements in deep learning architectures and datasets, reliable uncertainty estimation remains elusive, making CP increasingly vital. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning tasks, including classification, regression, graph neural networks, and large language models. TorchCP offers a comprehensive suite of advanced methodologies, a modular design for easy customization, and full GPU-accelerated scalability. Released under the LGPL-3.0 license, TorchCP has gained widespread adoption with over 12,582 PyPi downloads. It is supported by approximately 16,132 lines of code, 564 unit tests achieving 100% coverage, and comprehensive documentation. By bridging statistics and computer science, TorchCP empowers researchers and practitioners to advance conformal prediction in diverse deep learning applications.
@@ 1 Introduction
Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets that contain true labels with a specified coverage probability (Vovk et al., 2005). In recent years, deep learning has evolved dramatically with advanced model architectures and vast datasets, yet it struggles to reliably measure predictive uncertainty, making conformal prediction an increasingly attractive solution. Beyond classification and regression, conformal prediction methods have been tailored for diverse deep learning tasks, such as graph neural networks (GNNs) for molecular modeling and large language models (LLMs) for natural language understanding. This underscores the need for a comprehensive and efficient CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across diverse deep learning applications.
This work presents TorchCP, a PyTorch-native library of conformal prediction for deep learning. Firstly, TorchCP implements a wide range of state-of-the-art CP algorithms encompassing both training and inference methods designed for deep learning tasks, e.g., classification (ordinal classification), regression (time series), GNN, and LLM. Secondly, the library’s modular design allows developers to easily use or customize each component (e.g., losses or scores) of the library. Lastly, TorchCP is fully built on PyTorch with GPU support and batch processing, enabling seamless scalability for large-scale deep learning tasks.
The TorchCP library, released under the LGPL-3.0 license, can be installed via PyPi with the command ‘‘pip install torchcp’’ and has been downloaded 12,582 times  [^1] https://pepy.tech/projects/torchcp. The implementation consists of approximately 16,132 lines of code and is thoroughly tested with 564 unit tests, achieving 100% test coverage. We also provide comprehensive online documentation  [^2] https://torchcp.readthedocs.io/en/latest/ to guide users through TorchCP’s features. We hope the library can bridge the gap between statistics and computer science, enabling researchers and practitioners from both backgrounds to advance conformal prediction in deep learning.
2 The library design and implementation
Figure 1: An illustration of the modular design of TorchCP.
2.1 Modular design
For low coupling, we separate the TorchCP library into separate modules that can be easily maintained, tested, and integrated. For each task, the code implementation of the library contains 3 modules: Trainer, Score, and Predictor. Specifically, Trainer defines training algorithms for improving conformal prediction, along with loss functions. Score defines non-conformity score functions that quantify how different a new sample is from the training data, e.g., LAC (Sadinle et al., 2019). Predictor defines CP workflows (e.g., calibration and prediction), such as Split CP  [^3] A detailed introduction of Split CP is presented in Appendix A.1 (Papadopoulos et al., 2002). The modular design allows a straightforward extension to new tasks or new CP methods that arise in this active research field.
We visualize the modular design of TorchCP in Figure 1, highlighting the relationships among the three modules. The Trainer module is used to train or fine-tune deep models, employed by the Predictor module for calibration and prediction. The Score module implements non-conformity score functions, enabling their flexible use across different trainers and predictors. This modular structure empowers users to independently select or customize trainers, score functions, and predictors for diverse conformal prediction tasks, significantly improving flexibility and maintainability. Additionally, TorchCP achieves strong extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s modular design.
2.2 Implementation
@@
Table 1: The CP algorithms implemented in TorchCP.
Task Component Methods
@@
Classification Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
@@
Predictor Split CP (Papadopoulos et al., 2002) Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
Regression Score Function ABS (Papadopoulos et al., 2002), NORABS (Vovk et al., 2005), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
Predictor Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
@@
Table 1 outlines the CP algorithms included in TorchCP across 4 deep learning tasks: classification, regression (time series), GNN, and LLM , along with their corresponding evaluation metrics in Table 4. In total, TorchCP provides 6 training algorithms, 16 score functions, and 10 predictors tailored to various tasks. A brief overview of all supported methods and their advantages is available in Appendix A.
@@
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to real-world deep learning applications. Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
Tutorials
TorchCP aims to provide a user-friendly framework for conformal prediction in deep learning. The library contains illustrative code showcasing functionality for common tasks (See the “examples” folder  [^4] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples), and Appendix C offers a practical guide to help users identify relevant examples. In Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation. Below, we present a generic example  [^5] The full code of the example is available at https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_conftr_cifar100.py to show the usability of TorchCP. First, we employ ConfTr (Stutz et al., 2022) to train a model for conformal prediction (optional, using standard trained models also work):
@@
2trainer = ConfTrTrainer(init_model, alpha=0.1) # Create a trainer with a NN
@@
With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor. As an example, we implement Split CP with a miscoverage rate of $\alpha=0.1$ using the LAC score :
@@
2predictor.calibrate(cal_loader) # Compute the threshold on calibrate data
3predict_set = predictor.predict(test_instance) # Predict the set for an instance
@@ 3 Related library
Conformal Prediction (CP) is a distribution-free framework for quantifying the uncertainty of machine learning model predictions. While some uncertainty quantification libraries, such as PUNCC (Mendil et al., 2023), implement certain CP algorithms, their limited efficiency, scalability, and algorithmic coverage restrict their use in deep learning. For instance, most libraries support only a narrow range of tasks, such as regression and classification, and lack support for complex applications like generative modeling or time-series forecasting. Additionally, the absence of GPU acceleration and batch processing in these libraries makes them inefficient for large-scale scenarios. In contrast, TorchCP offers a clean interface, enhanced extensibility, high efficiency, and a comprehensive set of CP algorithms. For example, TorchCP reduces inference time by up to 90% compared to PUNCC on large-scale datasets (see Appendix D.2 for details). We summarize the comparison of different libraries in Table 2 and discuss the advantages of TorchCP ’s modular design in Appendix E.
Table 2: Feature comparison of open-source libraries supporting conformal prediction. “Training” refers to CP-specific training algorithms, and “Online” denotes online CP.
Library Task Support Functionality Support Backend
Classification Regression GNN LLM Training Online GPU Batch
nonconformist (Linusson et al., 2020) Numpy
TorchUQ (Zhao et al., 2024) PyTorch
Crepes (Boström, 2022) Numpy
PUNCC (Mendil et al., 2023) Numpy
MAPIE (Cordier et al., 2023) Numpy
Fortuna (Detommaso et al., 2024) JAX
TorchCP (ours) PyTorch
@@ 4 Conclusion and Future work
In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction (CP) into deep learning workflows. Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a modular design for easy customization and user-friendly documentation for accessibility. Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance. Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
@@ Appendix A Details of Conformal Prediction Methods in TorchCP
NORABS Improving robustness of prediction intervals via noise-aware normalization
@@ A.1 Split CP
Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth labels with a desired probability. Split Conformal Prediction (Papadopoulos et al., 2002) (Split CP) is the most widely-used version of the conformal prediction procedure. Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration. We next outline the main process of split conformal prediction:
@@
A.2 Classification
Here, we provide a brief overview of conformal prediction methods for multi-class classification tasks as supported by TorchCP. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
@@ A.2 Classification
$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\eta\cdot(r_{f}(\bm{x},y)-k_{reg})^{+},$
where $\eta$ represents the weight of regularization, $k_{reg}\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
@@
A.3 Regression
In the context of regression, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.

### v4

@@ Abstract
Conformal prediction (CP) is a powerful statistical framework that generates prediction intervals or sets with guaranteed coverage probability. While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale DL scenarios. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning techniques, including DNN-based classifier/regressor, GNN, and LLM. Released under the LGPL-3.0 license, TorchCP comprises about 16k lines of code, validated with 100% unit test coverage and detailed documentation. Notably, TorchCP enables CP-specific training algorithms, online prediction, and GPU-accelerated batch processing, achieving up to 90% reduction in inference time on large datasets. With its low-coupling design, comprehensive suite of advanced methods, and full GPU scalability, TorchCP empowers researchers and practitioners to enhance uncertainty quantification across cutting-edge applications.
@@ 1 Introduction
Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty. While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated DL models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability for adapting to large-scale DL scenarios. This underscores the need for a comprehensive and scalable CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
This work presents TorchCP, a PyTorch-native library that advances conformal prediction for deep learning by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifier/regressor, GNN, and LLM. Its low-coupling design facilitates easy customization of components like trainers or scores, while full GPU support and batch processing ensure seamless scalability for large-scale tasks. TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix D.2), while also encompassing a broader algorithmic coverage than TorchUQ. This PyTorch-native integration and extensibility empower statisticians and DL practitioners to seamlessly apply CP across cutting-edge applications with enhanced DL compatibility.
2 Main contributions
Conformal prediction constructs prediction sets $C(X)$ for input $X$, ensuring the true outcome $Y$ satisfies $\mathbb{P}(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate. Existing CP libraries encounter significant challenges in deep learning, including limited native support for modern models (like DNN, GNN, and LLM), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications. TorchCP overcomes these shortcomings through a PyTorch-native framework that improves model compatibility, efficiency, and adaptability, thereby enhancing CP’s applicability to DL. Its low-coupling design and GPU-accelerated batch processing provide a robust foundation for tackling complex, real-world problems, detailed in the following parts.
Figure 1: The modules of TorchCP.
2.1 Advantages for DL scenarios
Low-coupling design Low-coupling design is one of the most fundamental concepts in object-oriented design (McNatt and Bieman, 2001), which minimizes the dependency a class has on other classes. While many libraries (e.g., PUNCC) are implemented in a high-coupling design, TorchCP is structured into three core modules: Trainer, Score, and Predictor, each tailored to specific aspects of CP workflows. The Trainer module implements CP-specific training algorithms, such as ConfTr (Stutz et al., 2022), which optimizes deep models to improve efficiency. The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to training data. The Predictor module implements CP workflows, including calibration and prediction, to generate outputs, e.g., prediction sets/intervals, p-values, and conformal predictive distributions.
This low-coupling design is visualized in Figure 1, illustrating the interplay among modules. The Score module is used to create various score objects that serve as inputs for both the Trainer and Predictor modules. The Trainer module optimizes DL models using CP-specific loss functions. The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets, intervals, p-values, or predictive distributions. In this manner, TorchCP achieves strong reusability and extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s low-coupling design.
@@ 2.1 Advantages for DL scenarios
Table 1: The CP algorithms implemented in TorchCP for various deep learning models.
Model types Component CP Algorithms
@@
DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
@@
Predictor Split CP (Papadopoulos et al., 2002),Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
DNN Regressor Score Function ABS (Papadopoulos et al., 2002), NORABS (Papadopoulos et al., 2011), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
Predictor Split CP, Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
@@
Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifier/regressor, GNN, and LLM, along with their evaluation metrics in Table 4. Current libraries are typically limited in algorithmic coverage, e.g., MAPIE only includes split CP. In TorchCP, we provide 6 training algorithms, 17 scores, and 10 predictors. A brief overview of all supported methods and their advantages is available in Appendix A.
@@
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to deep learning applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)). Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
2.2 Tutorials
TorchCP provides a user-friendly framework for conformal prediction in deep learning with a detailed documentation [^2] https://torchcp.readthedocs.io/en/latest/index.html. The library contains illustrative code showcasing functionality for common applications (See the “examples” folder  [^3] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples) with a practical guide in Appendix C. Moreover, in Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation. Below, we present a generic example  [^4] The full source code is available at TorchCP CIFAR-100 Example. to show the usability of TorchCP. First, we utilize ConfTr (Stutz et al., 2022) to train a neural network, i.e., init_model, on a training dataset encapsulated in train_loader:
@@ 2.2 Tutorials
2trainer = ConfTrTrainer(init_model, alpha=0.1) # Initialize trainer with model
@@
With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor. Below, we implement Split CP with a miscoverage rate of $\alpha=0.1$, using the LAC score on calibration data encapsulated in cal_loader:
@@
2predictor.calibrate(cal_loader) # Set threshold with calibration data
3predict_set = predictor.predict(test_instance) # Predict set for test instance
@@ 3 Related library
Conformal prediction is a distribution-free framework for quantifying uncertainty, yet existing libraries face notable limitations in deep learning applications. Libraries such as nonconformist, PUNCC, and MAPIE offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets. Moreover, tools like TorchUQ and Fortuna provide some GPU support but fall short in algorithmic coverage, e.g., TorchUQ supports only regressors. Additionally, many libraries like PUNCC are implemented in a high-coupling design, where predictors and score functions are tightly coupled within a single class. TorchCP addresses these gaps with its PyTorch-native integration, supporting a broader range of modern models and offering a modular, extensible architecture, as detailed in Table 8. We present a detailed comparison of the relevant libraries in Appendix F.
Table 2: Comparison of Open-Source CP Libraries  [^7] The definition of model support and functionality support is presented in Appendix F.. “Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests but is primarily designed for advanced deep learning models within the PyTorch ecosystem.. The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction. “$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
Library Model Support Functionality Support Coupling Backend
Classifier Regressor GNN LLM Training Online GPU Batch $p$-value CPD
nonconformist (Linusson et al., 2020) low Numpy
TorchUQ (Zhao et al., 2024) low PyTorch
Crepes (Boström, 2022) low Numpy
PUNCC (Mendil et al., 2023) high Numpy
MAPIE (Cordier et al., 2023) high Numpy
Fortuna (Detommaso et al., 2024) high JAX
TorchCP (ours) low PyTorch
@@ 4 Conclusion and Future Work
In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction algorithms into deep learning workflows. Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a low-coupling design for easy customization and user-friendly documentation for accessibility. Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance. Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
@@ Appendix A Details of Conformal Prediction Methods in TorchCP
NORABS Enhancing Prediction Interval Robustness through Normalized Scoring
@@ A.1 Split CP
Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates prediction sets/intervals containing ground-truth labels with a desired probability. Split Conformal Prediction (Vovk et al., 2022) (Split CP) is the most widely-used version of the conformal prediction procedure. Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration. We next outline the main process of split conformal prediction:
@@
A.2 DNN Classifier
Here, we provide a brief overview of conformal prediction methods implemented using a DNN classifier with the TorchCP library. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a (KK-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
@@ A.2 DNN Classifier
$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\lambda\cdot(r_{f}(\bm{x},y)-k)^{+},$
where $\lambda$ represents the weight of regularization, $k\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
@@
A.3 DNN Regressor
For DNN Regressor, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.